### v1

# On continuity of state-dependent utilities

Abstract
State-dependent preferences for a general Savage’s state space were shown in Wakker and Zank (1999) to admit a numerical representation in the form of the integral of a state-dependent utility, as soon as pointwise continuity of the preference ordering is assumed. In this paper we prove that such a state-dependent function inherits pointwise continuity from the preference ordering, providing in this way a positive answer to a conjecture posed in the aforementioned seminal work.
We further apply this result to obtain an explicit representation of conditional Chisini means in the form of a conditional certainty equivalent.
keywords:
pointwise continuity, state-dependent expected utility, conditional certainty equivalent, nonlinear expectation
MSC:
[2020] 91B06, 91B08, 60A05
[…]
1 Introduction
For a general state space $(\Omega,{\mathcal{F}})$, Wakker and Zank WZ99 provide a characterization of those preference relations $\succeq$ on the space of bounded random acts which admit a numerical representation through an additively decomposable functional (WZ99 , Theorem 11). This result extends to an infinite dimensional setup the theory developed by Debreu De60 . By adopting the stronger notion of pointwise continuity of $\succeq$, Theorem 11 in WZ99 can be specialized to the following integral representation form.
Theorem 1.1 (WZ99 , Theorem 12 and CL06 , Theorem 5)
Assume that ${\mathcal{F}}$ contains at least three disjoint non-null events. Then the following two statements are equivalent:
  1. (i)
    There exists a countably additive finite measure $\mu$ on $\Omega$ and a function $\mathbb{u}(\omega,\cdot):\mathbb{R}\to\mathbb{R}$ strictly increasing for every $\omega\in\Omega$, such that $\succeq$ is represented by the pointwise continuous integral
    $f\to\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mu(\omega).$
  2. (ii)
    $\succeq$ is a preference order satisfying strict monotonicity, the sure-thing principle and pointwise continuity  [^1] See page (SM) below for further details on the preference axioms involved.
The previous theorem leaves an interesting issue unsolved as declared by the authors WZ99 , page 29:
“[…] Pointwise continuity is not overly restrictive because continuity of each state-dependent utility implies pointwise continuity. Whether the reversed implication holds, i.e., whether continuity of the functions $\mathbb{u}$, in Statement (i) of Theorem 12 holds (after appropriate modification on null events), is an open question to us. ”
To the best of our knowledge this question did not find an answer in the subsequent literature. Among the principal aims of the present paper, Theorem 3.1 stated below provides the construction of a pointwise continuous version of the state dependent utility $\mathbb{u}$. Such a result has the immediate benefit of clarifying the role of continuity in the application of the Theorem 1.1. For example, in Appendix A of MM21 a somewhat cumbersome notion of continuity is introduced to overcome some technical hurdles in the proof of the main result. Guaranteeing the existence of a continuous $\mathbb{u}$ allows to circumvent issues of this kind.
As an important byproduct, we provide a representation of the conditional Chisini mean $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ previously introduced and studied in DM23 . The conditional Chisini mean is a nonlinear extension of the standard concept of conditional expectation, which takes inspiration from the original Chisini’s intuition Ch29 that the notion of “expected value” should be introduced as the solution of a functional equation. For a sub-$\sigma$-algebra ${\mathcal{G}}$ of ${\mathcal{F}}$, representing intermediate information available, we show that $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)=\mathbb{u}_{\mathcal{G}}^{-1}(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right])$, where $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ is the conditional expectation of the random outcome $\omega\mapsto\mathbb{u}(\omega,f(\omega))$ and $\mathbb{u}_{\mathcal{G}}(\omega,x)=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(x)\middle|{\mathcal{G}}\right](\omega)$ is the ${\mathcal{G}}$-measurable projection of the state dependent utility $\mathbb{u}$. This representation can be easily compared to the general formulation proved in CVMM , Lemma 5.2, under additional law invariance assumptions, which traces back to the celebrated Nagumo, de Finetti, Kolmogorov Theorem (see HLP52 for a review). It is also worth mentioning that in this case, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ takes the form of a conditional certainty equivalent based on state-dependent utilities (see FM11cce ).
2 Notations and preliminaries
[…]
Whenever working with the product space $\Omega\times\mathbb{R}$, we shall make use of the product $\sigma$-algebra ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$ i.e. the $\sigma$-algebra generated by elements of the form $F\times B$ with $F\in{\mathcal{F}}$ and $B\in\mathbb{B}_{\mathbb{R}}$. Given a probability measure $\mathbb{P}$, $(\Omega,\mathcal{F},\mathbb{P})$ is a probability space. A (finite) signed measure $\nu$ on $(\Omega,{\mathcal{F}})$ is said to be dominated by $\mathbb{P}$ ($\nu\ll\mathbb{P}$) if, for any $A\in{\mathcal{F}}$, $\mathbb{P}(A)=0$ implies $\lvert\nu\rvert(A)=0$, where $\lvert\nu\rvert$ is the total variation measure of $\nu$. By the Hahn decomposition theorem, $\nu\ll\mathbb{P}$ is equivalent to: $\mathbb{P}(A)=0$ implies ${\nu}(A)=0$.
For a given probability $\mathbb{P}$ on $(\Omega,{\mathcal{F}})$ we shall denote the space of integrable functions by ${\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})=\left\{f\in{\mathcal{L}}(\Omega,{\mathcal{F}}):\mathbb{E}_{\mathbb{P}}\left[|f|\right]<\infty\right\}$.
For any $f\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ and a sub $\sigma$-algebra ${\mathcal{G}}$, $\mathbb{E}_{\mathbb{P}}\left[f\mid{\mathcal{G}}\right]$ represents the conditional expectation i.e.
$\mathbb{E}_{\mathbb{P}}\left[f\middle|{\mathcal{G}}\right]=\{g\in{\mathcal{L}}^{1}(\Omega,{\mathcal{G}},\mathbb{P}):\mathbb{E}_{\mathbb{P}}\left[g\mathbf{1}_{A}\right]=\mathbb{E}_{\mathbb{P}}\left[f\mathbf{1}_{A}\right]\;\forall\,A\in{\mathcal{G}}\}.$ (1)
In the following we shall assume familiarity with the properties of conditional expectation. We refer the reader to Williams , page 88, for a comprehensive review of such properties. A preference relation is a binary relation $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ which satisfies completeness and transitivity and induces a (strict) preference order $\succ$ by $f\succ g$ if and only if $f\succeq g$ but $g\nsucceq f$. In addition we set $g\sim f$ whenever $g\succeq f$ and $f\succeq g$. We denote
${\mathcal{N}}_{\succeq}:=\left\{A\in{\mathcal{F}}:f\mathbf{1}_{A}+g\mathbf{1}_{A^{c}}\sim g\;\;\forall f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})\right\}$ (2)
the set of null events (induced by the preference $\succeq$).
We present a few additional properties that the preference relation might enjoy.
$\succeq$ is strictly monotone if $x\mathbf{1}_{A}+f\mathbf{1}_{A^{c}}\succ y\mathbf{1}_{A}+f\mathbf{1}_{A^{c}}$, for all non null events $A\in{\mathcal{F}}$, for all $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and constant outcomes $x,y\in\mathbb{R}$ with $x>y$;
$\succeq$ satisfies the sure-thing principle whenever, considering arbitrary $f,g,h\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$ such that $f\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}\succeq g\mathbf{1}_{A}+h\mathbf{1}_{A^{c}}$, for every $\tilde{h}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we have $f\mathbf{1}_{A}+\tilde{h}\mathbf{1}_{A^{c}}\succeq g\mathbf{1}_{A}+\tilde{h}\mathbf{1}_{A^{c}}$;
$\succeq$ is pointwise continuous if, for any uniformly bounded sequence $(f_{n})_{n}\subseteq\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, such that $f_{n}(\omega)\rightarrow f(\omega)$ for any $\omega\in\Omega$ then for all $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ such that $g\succ f$ (resp. $f\succ g$) there exists $N\in\mathbb{N}$ such that $g\succ f_{n}$ (resp. $f_{n}\succ g$) for every $n>N$.
Assumptions (SM), (ST) and (PC) represent the minimal framework we will be working with throughout the paper. We recall that, provided ${\mathcal{F}}$ contains at least three distinct non-null events, by Theorem 1.1 any preference ordering satisfying these assumptions admits an integral representation.
3 Statement of the main results
We are now ready to state the two main results of this paper. The first one precisely answers to the question posed by WZ99 as announced in the introduction and from this perspective it provides an important addendum to Theorem 1.1.
Theorem 3.1
Assume that ${\mathcal{F}}$ contains at least three disjoint non-null events and $\succeq$ satisfies (SM), (PC) and (ST). Then there exist:
  1. (i)
    a probability measure $\mathbb{P}$ on ${\mathcal{F}}$,
  2. (ii)
    a ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ with $\mathbb{u}(\omega,\cdot)$ being a continuous and strictly increasing function on $\mathbb{R}$ for every $\omega\in\Omega$ and $\mathbb{u}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$,
such that given $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we have
$f\succeq g\text{ if and only if }\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\geq\int_{\Omega}\mathbb{u}(\omega,g(\omega))\mathrm{d}\mathbb{P}(\omega)$ (3)
In our second result we apply Theorem 3.1 to obtain an explicit formula for conditional Chisini means. Given a functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\rightarrow\mathbb{R}$ and a random variable $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, DM23 proposes sufficient conditions to guarantee the existence of a solution $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ to the infinite dimensional system of equations $T(f\mathbf{1}_{A})=T(g\mathbf{1}_{A})$ for any $A\in{\mathcal{G}}$. The family of solutions is the so-called conditional Chisini mean, denoted by $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$. If compared with (1), it evidently appears that $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ is the nonlinear counterpart of the standard conditional expectation. It is known that the problem of existence of nonlinear conditional expectation has been addressed also in CoquetPeng02 , yet the provided solution relies on a Brownian framework which in turn demands for more technical assumptions.
Embedding such a problem in a decision theoretical setup, we could rephrase the issue as follows: given a preference order $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ then the certainty equivalent of $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ is $a\in\mathbb{R}$ such that $a\sim f$. Similarly if the agent is provided with additional information, represented by the $\sigma$-algebra ${\mathcal{G}}\subset{\mathcal{F}}$, then the conditional certainty equivalent of $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ is any $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that
$f\mathbf{1}_{A}\sim g\mathbf{1}_{A}\quad\forall\;A\in{\mathcal{G}}.$ (4)
Loosely speaking, $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ can be considered as a ${\mathcal{G}}$-measurable counterpart of the ${\mathcal{F}}$-measurable function $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ (see also FM11cce and MM21 for applications of the conditional certainty equivalent).
Since this paper deals with state dependent preferences we shall directly refer to the conditional Chisini mean as the class
$\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)=\left\{g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})\mid f\mathbf{1}_{A}\sim g\mathbf{1}_{A}\;\forall\,A\in{\mathcal{G}}\right\},$ (5)
[…]
Theorem 3.2
[…]
  • 1.
    a function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$, with
    1. (i)
      $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ continuous and strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$,
    2. (ii)
      $\mathbb{u}_{\mathcal{G}}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{G}},\mathbb{P})$ for every $x\in\mathbb{R}$,
    3. (iii)
      $\mathbb{u}_{\mathcal{G}}(\cdot,x)$ is a version of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,x)\middle|{\mathcal{G}}\right]$ for every $x\in\mathbb{R}$.
  • 2.
    a version $g$ of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))\middle|{\mathcal{G}}\right]$ such that
    $\Phi_{\mathcal{G}}(\cdot,g(\cdot))\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right),$ (6)
where $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\rightarrow[-\infty,+\infty]$ defined by
$\Phi_{\mathcal{G}}(\omega,x):=\inf\{y\in\mathbb{R}\mid\mathbb{u}_{\mathcal{G}}(\omega,y)>x\}$
is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable.
We write informally (6) as
$\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)=\mathbb{u}_{\mathcal{G}}^{-1}(\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]),$
and recall that, under the assumptions of Theorem 3.2, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ is unique up to $\succeq$-null events (see DM23 , Theorem 3.2). To clarify the statement of Theorem 3.2, it seems important at this stage to point out that the use of $\Phi_{\mathcal{G}}$ in (6) conceals some technical subtleties. To begin with, the dependence of the image of $\mathbb{u}_{\mathcal{G}}$ on $\omega$ requires cautious handling of its domain. Additionally, for $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, the nature of $g\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ implies that it may not be an element of $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ itself. While the case of a finitely generated $\sigma$-algebra is rather simple (as we shall depict in Example 3.3 below), we shall dedicate a comprehensive treatment to the issue for an arbitrary $\sigma$-algebra ${\mathcal{G}}$ in the proof postponed in Section 5.
Example 3.3
Consider $\{A_{1},...,A_{n}\}$ being a finite partition of $\Omega$ such that $\mathbb{P}(A_{i})>0$ for every $i\leq n$ and the $\sigma$-algebra ${\mathcal{G}}$ generated by this partition. Then every ${\mathcal{G}}$-measurable function $g$ can be written as a finite linear combination of indicator functions of the atoms $A_{i}$. In particular, for $\omega\in A_{i}$, $\mathbb{u}_{\mathcal{G}}(\omega,x)$ coincides with
$u_{i}(x)=\frac{1}{\mathbb{P}(A_{i})}\int_{A_{i}}\mathbb{u}(\omega,x)\mathrm{d}\mathbb{P}(\omega),$
$u_{i}$ being finite since $\mathbb{u}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$. Clearly $\mathbb{u}_{\mathcal{G}}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{G}},\mathbb{P})$ for all $x\in\mathbb{R}$, $\mathbb{u}_{\mathcal{G}}(\omega,x)=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,x)|{\mathcal{G}}\right](\omega)$ and $x\mapsto\mathbb{u}_{\mathcal{G}}(\omega,x)$ is increasing and continuous for all $\omega\in\Omega$.
For any fixed $A_{i}$ we notice that the image of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$, namely $R_{i}:=\mathrm{Im}(\mathbb{u}_{\mathcal{G}}(\omega,\cdot))$ does not depend on $\omega\in A_{i}$. Given the generalized inverse $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ defined in Theorem 3.2, it is crucial to observe that the map $x\mapsto\Phi_{\mathcal{G}}(\omega,x)=\mathbb{u}_{\mathcal{G}}^{-1}(\omega,x)$ is well defined as a standard inverse function for $\omega\in A_{i}$ and $x\in R_{i}$ (as a result of strict monotonicity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$). In particular if we now choose any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and set $h:=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right]$ then $h(\omega)\in R_{i}$ for $\omega\in A_{i}$ and indeed
$(\mathbb{u}_{\mathcal{G}}\circ\Phi_{\mathcal{G}})(\omega,h(\omega))=h(\omega).$
For $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ denote $\mathfrak{m}_{i}=u_{i}^{-1}\left(\frac{1}{\mathbb{P}(A_{i})}\int_{A_{i}}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\right)$ so that for any $\omega\in A_{i}$ we have
$\mathbb{u}_{\mathcal{G}}(\omega,\mathfrak{m}_{i})=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right](\omega).$
We conclude the example noting that the Chisini mean $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ for the functional $T(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ consists, in the current setup, of a single random variable $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ which assumes the value
$u_{i}^{-1}\left(\frac{1}{\mathbb{P}(A_{i})}\int_{A_{i}}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\right)$
on the set $A_{i}$. This can be readily verified following the procedure in Section 5, step 4.
4 Proof of Theorem 3.1
This section is devoted to the proof of Theorem 3.1, which is quite involved and requires some preliminary groundwork.
In the first part of the proof we lay the basis for the main argument by rephrasing some known results from WZ99 ; DM23 , obtaining in particular a functional that represents $\succeq$, which turns out to be additive on measurable sets. We provide some auxiliary lemmata: Lemma 4.6 proves that it is possible to construct an integrand function with the desired properties; this is based on Lemma 4.4 which firstly ensures the suitable measurability by a construction on rational numbers; Lemma 4.7 states that the integral representation is well defined and enjoys the desired properties; finally Lemma 4.8 shows continuity of $\mathbb{u}$. The concluding argument leads to Proposition 4.1 and illustrates how ultimately Theorem 3.1 follows. Consider a preference order $\succeq$ on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ satisfying (SM), (PC) and (ST). Suppose additionally that ${\mathcal{F}}$ contains at least three disjoint events $A_{1},A_{2},A_{3}$ such that $A_{i}\notin{\mathcal{N}}_{\succeq}$ for $i=1,2,3$. It is known (see WZ99 , Theorem 11) that $\succeq$ admits a numerical representation, i.e. there exists a functional $T:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ such that for any given $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$
$T(f)\geq T(g)\quad\text{ if and only if }\quad f\succeq g.$
In particular WZ99 , Theorem 11, or CL06 , Theorem 4, imply that the Assumptions in DM23 , Theorem 5.4333Adopting ${\mathcal{F}}$ in place of ${\mathcal{G}}$, are met by the functional $T$ and we can therefore provide a further refinement of the representing functional. Specifically one can find a $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R},\;(A,f)\mapsto V_{A}(f)$ such that for $f_{1},f_{2}\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$ we have $f_{2}\mathbf{1}_{A}\succeq f_{1}\mathbf{1}_{A}$ if and only if $V_{A}(f_{1})\leq V_{A}(f_{2})$. Furthermore, $V$ can be taken to satisfy the following properties (we refer the reader to the aforementioned literature for further details):
  1. (P1)
    for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ the map $A\mapsto V_{A}(f)$ is a signed measure on $(\Omega,{\mathcal{F}})$ with $V_{A}(0)=0$ for every $A\in{\mathcal{F}}$. Moreover, $A\mapsto V_{A}(\mathbf{1}_{\Omega})=:\mathbb{P}(A)$ is a probability measure on $(\Omega,{\mathcal{F}})$;
  2. (P2)
    for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $A\in{\mathcal{F}}$ we have $V_{A}(f)=V_{\Omega}(f\mathbf{1}_{A})$;
  3. (P3)
    the functional $V$ is strictly monotone in that for any $f,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ such that $f\leq g$ then $V_{A}(f)\leq V_{A}(g)$ for any $A\in{\mathcal{F}}$, and if $A$ is such that $\mathbb{P}(A)>0$ then $V_{A}(x)<V_{A}(y)$ for any $x,y\in\mathbb{R}$ with $x<y$;
  4. (P4)
    for every $A\in{\mathcal{F}}$ the functional $f\mapsto V_{A}(f)$, defined on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, is pointwise continuous, in that for any uniformly bounded [^4] Meaning $\sup_{n}\lVert f_{n}\rVert_{\infty}<+\infty$ sequence $(f_{n})_{n}\subset\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $f_{n}(\omega)\to f(\omega)$ for any $\omega\in\Omega$, then $V_{A}(f_{n})\to V_{A}(f)$.
The following proposition provides a general representation result which stretches somewhat beyond the scope of the present work. A similar problem is in fact addressed within some relevant contributions in the stream of literature of AZ90 and (Buttazzo, , Ch. 2), under the name of Nonlinear Superposition Operator and exploiting the theory of Carathéodory functions. Another application of these representation results to subjective expected utility is presented in STANCA20 .
Proposition 4.1
[…]
$V_{A}(f)=\int_{A}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\quad\text{for any }f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F}),A\in{\mathcal{F}}.$
Remark 4.2
We emphasize that Proposition 4.1 provides a more general result than Theorem 3.1. Intuitively speaking, a functional that represents a preference relation as in WZ99 always satisfies the assumptions of the proposition, as will be explained more in detail at the end of this section.
Consider a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ satisfying (P1), (P2), (P3) and (P4). For any $A\in{\mathcal{F}},f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we denote
$\mu_{f}(A):=V_{\Omega\cap A}(f)=V_{A}(f),$
and observe that, for all $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $\mu_{f}$ is dominated by the probability $\mathbb{P}$ defined as in (P1). We now construct a special version of the state dependent utility, which enjoys the desired properties.
Definition 4.3
For any $q\in{\mathbb{Q}}$, we let $u(\cdot,q)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ be a version of the Radon-Nikodym derivative $\frac{\mathrm{d}\mu_{q}}{\mathrm{d}\mathbb{P}}$. In particular, we set $u(\omega,0)=0$ for all $\omega\in\Omega$ by virtue of (P1).
For any $A\in{\mathcal{F}}$ and $q\in\mathbb{Q}$, Definition 4.3 motivates the following notation
$V_{A}(q)=\int_{A}\mathrm{d}\mu_{q}(\omega)=\int_{\Omega}u(\omega,q)\mathbf{1}_{A}(\omega)\mathrm{d}\mathbb{P}(\omega)=\int_{\Omega}u(\omega,q\mathbf{1}_{A}(\omega))\mathrm{d}\mathbb{P}(\omega)$ (7)
Lemma 4.4
Denote by
$A^{\mathbb{Q}}:=\bigl{\{}\omega\in\Omega:u(\omega,q_{1})<u(\omega,q_{2}),\;\forall\;q_{1}<q_{2}\in{\mathbb{Q}}\bigr{\}}$
and
$B^{\mathbb{Q}}:=\bigl{\{}\omega\in\Omega:u(\omega,q)=\inf_{{\mathbb{Q}}\ni\tilde{q}>q}u(\omega,\tilde{q}),\;\forall\;q\in{\mathbb{Q}}\bigl{\}}.$
Then, $A^{\mathbb{Q}},B^{\mathbb{Q}}\in{\mathcal{F}}$ and $\mathbb{P}(A^{\mathbb{Q}})=\mathbb{P}(B^{\mathbb{Q}})=1$.
Proof 1
We start fixing $q_{1}<q_{2}\in{\mathbb{Q}}$ and denoting by $C=\{\omega\in\Omega\mid u(\omega,q_{1})\geq u(\omega,q_{2})\}$. By (7) we can write
$V_{C}(q_{1})=\int_{C}u(\omega,q_{1})\mathrm{d}\mathbb{P}(\omega)\geq\int_{C}u(\omega,q_{2})\mathrm{d}\mathbb{P}(\omega)=V_{C}(q_{2}),$
which contradicts (P3) unless $\mathbb{P}(C)=0$. We conclude that ($\mathbb{P}(\{\omega\in\Omega:u(\omega,q_{1})<u(\omega,q_{2})\})=1$. By (countable) intersection over the rational numbers we obtain
$A^{\mathbb{Q}}=\bigcap_{\begin{subarray}{c}q_{1},q_{2}\in{\mathbb{Q}},\\ q_{1}<q_{2}\end{subarray}}\bigl{\{}\omega\in\Omega:u(\omega,q_{1})<u(\omega,q_{2})\bigr{\}}$
and clearly $A^{\mathbb{Q}}\in{\mathcal{F}}$, with $\mathbb{P}(A^{\mathbb{Q}})=1$.
We now turn our attention to $B^{\mathbb{Q}}$. We show that
($\mathbb{P}(\{\omega\in\Omega:u(\omega,q)\leq\inf_{\tilde{q}>q}u(\omega,\tilde{q}),\;\forall\;q\in{\mathbb{Q}}\})=1$ (8)
and that
($\mathbb{P}(\{\omega\in\Omega:u(\omega,q)\geq\inf_{\tilde{q}>q}u(\omega,\tilde{q}),\;\forall\;q\in{\mathbb{Q}}\})=1,$ (9)
so that (8) and (9) together imply our claim. First of all we show that (8) holds. Fix $q,\tilde{q}\in{\mathbb{Q}}$ such that $\tilde{q}>q$, the previous argument shows that the event $\{\omega\in\Omega:u(\omega,q)\leq u(\omega,\tilde{q})\}$ has probability $1$. Notice that for $q\in{\mathbb{Q}}$ fixed, we have
$\bigcap_{\tilde{q}>q}\{\omega\in\Omega:u(\omega,q)\leq u(\omega,\tilde{q})\}=\{\omega\in\Omega:u(\omega,q)\leq\inf_{\tilde{q}>q}u(\omega,\tilde{q})\}.$
We denote by $E_{q}$ the event $\bigcap_{\tilde{q}>q}\{\omega\in\Omega:u(\omega,q)\leq u(\omega,\tilde{q})\}$. Naturally, the countable intersection is still measurable and in particular $\mathbb{P}(E_{q})=1$. Finally, since
$\bigcap_{q\in{\mathbb{Q}}}E_{q}=\{\omega\in\Omega:u(\omega,q)\leq\inf_{\tilde{q}>q}u(\omega,\tilde{q}),\;\forall\;q\in{\mathbb{Q}}\},$
we find $\mathbb{P}\left(\cap_{q\in{\mathbb{Q}}}E_{q}\right)=1$ and (8) holds.
We proceed proving (9) by contradiction and set
$A=\{\omega\in\Omega:u(\omega,q)<\inf_{\tilde{q}>q}u(\omega,\tilde{q})\}.$
Suppose that $\mathbb{P}\left(A\right)>0$, then, for some $n\in\mathbb{N}$, $A_{n}:=\{\omega\in\Omega:u(\omega,q)+\frac{1}{n}<\inf_{\tilde{q}>q}u(\omega,\tilde{q})\}$ is such that $\mathbb{P}(A_{n})>0$. Then it holds that
$\int_{A_{n}}\left(u(\omega,q)+\frac{1}{n}\right)\mathrm{d}\mathbb{P}(\omega)\leq\int_{A_{n}}\inf_{\tilde{q}>q}u(\omega,\tilde{q})\mathrm{d}\mathbb{P}(\omega)\leq\inf_{\tilde{q}>q}\int_{A_{n}}u(\omega,\tilde{q})\mathrm{d}\mathbb{P}(\omega).$ (10)
It follows from (10) that
$V_{A_{n}}(q)+\frac{1}{n}\mathbb{P}(A_{n})\leq\inf_{\tilde{q}>q}V_{A_{n}}(\tilde{q}).$ (11)
Since by (P3) and (P4) $V$ is strictly monotone and pointwise continuous we have that $\inf_{\tilde{q}>q}V_{A_{n}}(\tilde{q})=V_{A_{n}}(q)$, which in turn yields the desired contradiction. To conclude, since both (8) and (9) hold with probability $1$, we have
$\mathbb{P}(B^{\mathbb{Q}})=\mathbb{P}(\{\omega\in\Omega:u(\omega,q)=\inf_{\tilde{q}>q}u(\omega,\tilde{q}),\;\forall\;q\in{\mathbb{Q}}\})=1.$
We define the set
$\Theta=A^{\mathbb{Q}}\cap B^{\mathbb{Q}}\in{\mathcal{F}}\text{ with }\mathbb{P}(\Theta)=1,$ (12)
and recall that for any $\omega\in\Theta$ we have $u(\omega,q_{1})<u(\omega,q_{2})$ for all $q_{1}<q_{2}$, $q_{1},q_{2}\in{\mathbb{Q}}$ and $u(\omega,q)=\inf_{\tilde{q}>q}u(\omega,\tilde{q})$ for every $q\in{\mathbb{Q}}$.
Definition 4.5
For any $x\in\mathbb{R}$ and $\omega\in\Omega$ we define
$u^{+}(\omega,x)=\inf_{\begin{subarray}{c}q\in{\mathbb{Q}},\\ q\geq x\end{subarray}}\{u(\omega,q)\mathbf{1}_{\Theta}(\omega)+x\mathbf{1}_{\Omega\setminus\Theta}(\omega)\}.$
We observe that $u^{+}(\omega,x)$ is ${\mathcal{F}}$-measurable for any $x\in\mathbb{R}$, as it is the countable pointwise infimum of measurable functions. Notice that $u^{+}(\omega,0)=0$ for all $\omega\in\Omega$.
Lemma 4.6
Let $u^{+}:\Omega\times\mathbb{R}\to\mathbb{R}$ be as in Definition 4.5. Then the following hold:
  1. (i)
    $u^{+}(\omega,q)=u(\omega,q)$ for any $\omega\in\Theta$ (as defined in (12)) and $q\in{\mathbb{Q}}$;
  2. (ii)
    if $x<y$, $u^{+}(\omega,x)<u^{+}(\omega,y)$ for every $\omega\in\Omega$;
  3. (iii)
    $u^{+}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$;
  4. (iv)
    for every $\omega\in\Omega$ the function $u^{+}(\omega,\cdot)$ is RCLL555namely right continuous with left limits on $\mathbb{R}$;
Proof 2
Item (i) follows from the definition of $\Theta$. Item (ii) holds trivially on $\Omega\setminus\Theta$. Take now any $\omega\in\Theta$: provided that $x<y$, there exist $q_{1},q_{2}\in{\mathbb{Q}}$ such that $x\leq q_{1}<q_{2}\leq y$. By definition of $\Theta$ we have $u(\omega,q_{1})<u(\omega,q_{2})$ and moreover by monotonicity of $u$ and definition of $u^{+}$
$u(\omega,q)\leq u^{+}(\omega,y),\quad\text{for every }q\in{\mathbb{Q}},y\in\mathbb{R}\text{ s.t. }q\leq y.$
Then, we see
$u^{+}(\omega,x)=\inf_{\begin{subarray}{c}q\in{\mathbb{Q}},\\ q\geq x\end{subarray}}u(\omega,q)\leq u(\omega,q_{1})<u(\omega,q_{2})\leq\inf_{\begin{subarray}{c}q\in{\mathbb{Q}},\\ q\geq y\end{subarray}}u(\omega,q)=u^{+}(\omega,y).$
As to item (iii), fix any $x\in\mathbb{R}$ and consider $q_{1},q_{2}\in{\mathbb{Q}}$ such that $q_{1}\leq x\leq q_{2}$. Then by definition and item (i) $u(\omega,q_{1})\leq u^{+}(\omega,x)\leq u(\omega,q_{2})$ for any $\omega\in\Theta$. Since $\mathbb{P}(\Theta)=1$ and $u(\cdot,q_{1}),u(\cdot,q_{2})$ are versions of the Radon-Nikodym derivatives $\frac{d\mu_{q_{1}}}{\mathrm{d}\mathbb{P}},\frac{d\mu_{q_{2}}}{\mathrm{d}\mathbb{P}}\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$, integrability follows. Finally, observe that item (iv) holds trivially for $\omega\in\Omega\backslash\Theta$. Fix now $\omega\in\Theta$, we wish to show that for any $\varepsilon>0$ there exists $\delta$ such that $|u^{+}(\omega,x)-u^{+}(\omega,y)|<\varepsilon$ for any $y\in(x,x+\delta)$. Fix $\varepsilon>0$ and let $(q_{n})_{n}\subset{\mathbb{Q}}$ be such that $q_{n}\downarrow_{n}x$. Then there exists $\bar{n}\in\mathbb{N}$ such that, for all $n>\bar{n}$ we have $|u(\omega,q_{n})-\inf_{q>x}u(\omega,q)|=u(\omega,q_{n})-\inf_{q>x}u(\omega,q)<\varepsilon$. Recalling that, from item (i), $u(\omega,q_{n})=u^{+}(\omega,q_{n})$ and $\inf_{q>x}u(\omega,q)=u^{+}(\omega,x)$ by definition, monotonicity of $u^{+}$ implies that any $y\in(x,q_{\bar{n}})\subset\mathbb{R}$ satisfies $|u^{+}(\omega,x)-u^{+}(\omega,y)|<\varepsilon$. Existence of the left limit follows again from monotonicity proved in item (ii).
Lemma 4.7
Let $u^{+}:\Omega\times\mathbb{R}\to\mathbb{R}$ be as in Definition 4.5. Then the map
$f\mapsto T_{u^{+}}(f):=\int_{\Omega}u^{+}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ (13)
is well defined, monotone with respect to the pointwise order in $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, pointwise continuous [^6] As previously mentioned: for any uniformly bounded sequence $(f_{n})_{n}\subset\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $f_{n}(\omega)\to f(\omega)$ for any $\omega\in\Omega$, then $T_{u^{+}}(f_{n})\to T_{u^{+}}(f)$ and it satisfies $V_{A}(f)=T_{u^{+}}(f\mathbf{1}_{A})$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and $A\in{\mathcal{F}}$.
Proof 3
Observe that, since $u^{+}(\omega,\cdot)$ is RCLL for every $\omega\in\Omega$, by Lemma A.1 we have that $T_{u^{+}}$ is well defined on $\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ and monotone as $f(\omega)\leq g(\omega)$ implies $u^{+}(\omega,f(\omega))\leq u^{+}(\omega,g(\omega))$. We now show that $V_{\Omega}(f)=T_{u^{+}}(f)$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, then pointwise continuity of $T_{u^{+}}$ will follow from (P4). Fix any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, we can find two sequences of simple functions $\underline{f_{n}}=\sum_{A\in\underline{\pi}_{n}}\underline{x}^{n}_{A}\mathbf{1}_{A}$, $\overline{f_{n}}=\sum_{B\in\overline{\pi}_{n}}\overline{x}^{n}_{B}\mathbf{1}_{B}$, with $\overline{\pi}_{n},\underline{\pi}_{n}\subseteq{\mathcal{F}}$ finite partitions of $\Omega$, such that $\underline{f_{n}}\uparrow_{n}f,\overline{f_{n}}\downarrow_{n}f$. Moreover by standard arguments we can assume for any $n$ that $\underline{x}^{n}_{A},\overline{x}^{n}_{B}\in\mathbb{Q}$ for any $A\in\underline{\pi}_{n}$ and $B\in\overline{\pi}_{n}$. We have as a consequence of (7) and Lemma 4.6
$V_{\Omega}(\underline{f}_{n})$ $=\sum_{A\in\underline{\pi}_{n}}\int_{\Omega}u(\omega,\underline{x}_{n}^{A})\mathbf{1}_{A}\mathrm{d}\mathbb{P}(\omega)=\sum_{A\in\underline{\pi}_{n}}\int_{\Omega}u^{+}(\omega,\underline{x}_{n}^{A})\mathbf{1}_{A}\mathrm{d}\mathbb{P}(\omega)$
$=\int_{\Omega}u^{+}(\omega,\underline{f}_{n}(\omega))\mathrm{d}\mathbb{P}(\omega)\leq\int_{\Omega}u^{+}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$
and by a similar argument we can get
$\int_{\Omega}u^{+}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\leq\int_{\Omega}u(\omega,\overline{f}_{n}(\omega))\mathrm{d}\mathbb{P}(\omega)=V_{\Omega}(\overline{f}_{n})$
By (P4) we get the desired equality $V_{\Omega}(f)=T_{u^{+}}(f)$ for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$. Finally, we see that (P2) yields $V_{A}(f)=V_{\Omega}(f\mathbf{1}_{A})=T_{u^{+}}(f\mathbf{1}_{A})$, which concludes the proof.
Lemma 4.8
There exists a function $\mathbb{u}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that $\mathbb{u}(\omega,\cdot)$ is continuous and strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$, $\mathbb{u}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$ and
$V_{A}(f)=\int_{A}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)\quad\forall\;f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})\text{ and }A\in{\mathcal{F}}.$ (14)
Proof 4
We observe that, by Lemma 4.6, $u^{+}(\omega,\cdot)$ is strictly increasing, RCLL on $\mathbb{R}$ for all $\omega\in\Omega$ and $u^{+}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$. Moreover, as argued in Lemma 4.7, $T_{u^{+}}$ is pointwise continuous and in particular continuous from below. Consequently the assumptions of Corollary A.3 are verified and its application yields the function $\widehat{u^{+}}$ which we rename $\mathbb{u}:=\widehat{u^{+}}$. Finally, $\mathbb{u}$ satisfies all the requirements since
[…]
Conclusions. We conclude the section by proving Proposition 4.1 and we show that Theorem 3.1 follows. Let $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ be a functional satisfying (P1), (P2), (P3) and (P4). Lemma 4.4 and 4.6 provide the construction of a suitable integrand function $u^{+}$ and Lemma 4.7 ensures that the map $f\mapsto\int_{\Omega}u^{+}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ is well defined. Lemma 4.8 concludes the proof by showing the existence of the desired version $\mathbb{u}$ with the required properties. To prove Theorem 3.1, we exploit Proposition 4.1. Consider a preference $\succeq$ that is (SM), (PC) and (ST) and assume that ${\mathcal{F}}$ contains at least three disjoint events $A_{1},A_{2},A_{3}$ such that $A_{i}\notin{\mathcal{N}}_{\succeq}$ for $i=1,2,3$. As argued in the beginning of Section 4, the preference relation $\succeq$ induces a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ that satisfies (P1), (P2), (P3) and (P4)  with $V_{\Omega}(\cdot)$ acting as numerical representation of the preference $\succeq$. Application of Proposition 4.1 provides a pair $(\mathbb{u},\mathbb{P})$ such that $V_{\Omega}(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$.
5 Proof of Theorem 3.2
We begin by recalling that we work under the hypotheses of Theorem 3.1. In analogy to the previous section, it is possible to construct a functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ that satisfies (P1), (P2), (P3) and (P4). We also point out that properties (P1-4) are in fact stronger than those required in DM23 , Theorem 3.2777Indeed, using now the properties defined in DM23 , the fact that $V_{\Omega}:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ is (${\mathcal{G}}$-Mo), (${\mathcal{G}}$-PC) and (${\mathcal{G}}$-NB) follows trivially from (P3) and (P4), for any sub-$\sigma$-algebra ${\mathcal{G}}$ of ${\mathcal{F}}$. Furthermore, recalling that, by (P1), $V_{A}(0)=0$ we are able to recover also (${\mathcal{G}}$-QL) and (${\mathcal{G}}$-PS). Notice that, for $g_{1},g_{2}\in{\mathcal{L}}^{\infty}(\Omega,{\mathcal{G}}),A\in{\mathcal{G}}$, it is possible to write $V_{\Omega}(g_{1}\mathbf{1}_{A}+g_{2}\mathbf{1}_{A^{c}})=V_{A}(g_{1})+V_{A^{c}}(g_{2})$.. Therefore, $\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)\neq\emptyset$ and for any $g_{1},g_{2}\in\mathfrak{m}\left(f\middle|{\mathcal{G}}\right)$ we have $\{g_{1}\neq g_{2}\}\in{\mathcal{N}}_{\succeq}$. Step 1: We first show the existence of a state dependent utility on $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ for any $\sigma$-algebra ${\mathcal{G}}\subseteq{\mathcal{F}}$. Consider the functional $V:{\mathcal{F}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{F})\to\mathbb{R}$ defined above. By Proposition 4.1 it follows that $V$ admits a representation of the kind $V_{A}(f)=\int_{A}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$. Let ${\mathcal{G}}\subseteq{\mathcal{F}}$ be an arbitrary sub-$\sigma$-algebra on $\Omega$ and denote by $V^{\mathcal{G}}:{\mathcal{G}}\times\mathcal{L}^{\infty}(\Omega,\mathcal{G})\to\mathbb{R}$ the restriction of the functional $V$ to ${\mathcal{G}}\subseteq{\mathcal{F}}$ and $\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. We stress that, differently from ${\mathcal{F}}$, the $\sigma$-algebra ${\mathcal{G}}$ need not contain at least three disjoint events of positive probability. Analogously to $V$, $V^{\mathcal{G}}$ defines a probability measure $\mathbb{P}_{\mathcal{G}}$ on ${\mathcal{G}}$, which is the restriction of the probability measure $\mathbb{P}$ induced by $V$ to the sub-$\sigma$-algebra ${\mathcal{G}}$. We will hence use only $\mathbb{P}$ in the following, to avoid impractical notations.
Applying Proposition 4.1 to $V^{\mathcal{G}}$ it is possible to construct a ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$-measurable function $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ such that
  1. (i)
    $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ is continuous and strictly increasing on $\mathbb{R}$ for all $\omega\in\Omega$,
  2. (ii)
    $\mathbb{u}_{\mathcal{G}}(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{G}},\mathbb{P})$ for all $x\in\mathbb{R}$,
  3. (iii)
    $V_{A}^{\mathcal{G}}(g)=\int_{A}\mathbb{u}_{\mathcal{G}}(\omega,g(\omega))\mathrm{d}\mathbb{P}(\omega),\,\forall A\in{\mathcal{G}},g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$.
Notice that for all $A\in{\mathcal{G}}$ and $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$
$V_{A}^{\mathcal{G}}(g)=\int_{A}\mathbb{u}_{\mathcal{G}}(\omega,g(\omega))\mathrm{d}\mathbb{P}(\omega)=\int_{A}\mathbb{u}(\omega,g(\omega))\mathrm{d}\mathbb{P}(\omega)=V_{A}(g).$
It is then immediate to verify by definition of conditional expectation that the random variable $\mathbb{u}_{\mathcal{G}}(\cdot,x)$ is an element of the class $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,x)|{\mathcal{G}}\right]$. To avoid overburdening notation, in the following step 2 we shall write $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)\middle|{\mathcal{G}}\right]$ in place of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f(\cdot))|{\mathcal{G}}\right]$ and $\mathbb{u}_{\mathcal{G}}(g)$ instead of $\mathbb{u}_{\mathcal{G}}(\cdot,g(\cdot))$. Moreover, we shall often make the following standard abuse of notation: for any $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$ we write $g=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$ to indicate that $g\in\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$ i.e. $g$ is a version of the conditional expectation $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(f)|{\mathcal{G}}\right]$. Step 2: We show that:
$\mathbb{u}_{\mathcal{G}}(g)\text{ belongs to }\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(g)\middle|{\mathcal{G}}\right]\quad\forall\,g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G}).$ (15)
Consider indeed a simple function $g=\sum_{A\in\pi}y_{A}\mathbf{1}_{A}$, with $\pi$ being a finite ${\mathcal{G}}$-measurable partition of $\Omega$. Then we have
$\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(g)\middle|{\mathcal{G}}\right]$ $=\sum_{A\in\pi}\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(g)\mathbf{1}_{A}\middle|{\mathcal{G}}\right]=\sum_{A\in\pi}\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(y_{A})\mathbf{1}_{A}\middle|{\mathcal{G}}\right]$
$=\sum_{A\in\pi}\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(y_{A})\middle|{\mathcal{G}}\right]\mathbf{1}_{A}=\sum_{A\in\pi}\mathbb{u}_{\mathcal{G}}(y_{A})\mathbf{1}_{A}=\mathbb{u}_{\mathcal{G}}(\sum_{A\in\pi}y_{A}\mathbf{1}_{A}),$
and therefore (15) is verified for ${\mathcal{G}}$-measurable simple functions. We proceed to show it also holds for general functions $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. Take any sequence of ${\mathcal{G}}$-measurable simple functions $(\underline{g}_{n})_{n}\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$ such that $\underline{g}_{n}\uparrow_{n}g$ (again, the existence of such functions is guaranteed by standard arguments). Monotonicity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ ensures that $(\mathbb{u}_{\mathcal{G}}(\omega,\underline{g}_{n}(\omega)))_{n}$ is an increasing sequence for all $\omega\in\Omega$. Finally, pointwise continuity of $\mathbb{u}$ gives
$\mathbb{u}_{\mathcal{G}}(g)=\lim_{n}\mathbb{u}_{\mathcal{G}}(\underline{g}_{n}),$
and monotone convergence theorem yields
$\lim_{n}\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\underline{g}_{n})\middle|{\mathcal{G}}\right]=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(g)\middle|{\mathcal{G}}\right]$
which provides (15) for general $g\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. Step 3: we now study the generalized inverse of the state dependent utility $\mathbb{u}_{\mathcal{G}}$. If on the one hand $(\omega,x)\mapsto\mathbb{u}_{\mathcal{G}}(\omega,x)$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable (by Proposition 4.1), we want the generalized inverse $\Phi_{\mathcal{G}}$ to enjoy the same measurability as $\mathbb{u}_{\mathcal{G}}$. We first extend the domain of $\mathbb{u}_{\mathcal{G}}:\Omega\times\mathbb{R}\to\mathbb{R}$ to $\Omega\times\overline{\mathbb{R}}$ by denoting
$\mathbb{u}_{\mathcal{G}}(\omega,+\infty):=\lim_{x\to+\infty}\mathbb{u}_{\mathcal{G}}(\omega,x),\quad\quad\mathbb{u}_{\mathcal{G}}(\omega,-\infty):=\lim_{x\to-\infty}\mathbb{u}_{\mathcal{G}}(\omega,x).$
Monotonicity and continuity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ for all $\omega\in\Omega$ ensure that the limits exist. Fixed $\omega\in\Omega$, we denote with $\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}}):=\text{Im}(\mathbb{u}_{\mathcal{G}}(\omega,\cdot))$ the image of $\mathbb{u}_{\mathcal{G}}$, that is the interval $(\mathbb{u}_{\mathcal{G}}(\omega,-\infty),\mathbb{u}_{\mathcal{G}}(\omega,+\infty))$, and we stress that it is $\omega$-dependent and open by strict monotonicity.
Let $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\overline{\mathbb{R}}$ be the $\omega$-wise right-continuous generalized inverse:
$\Phi_{\mathcal{G}}(\omega,x):=\inf\{y\in\mathbb{R}\mid\mathbb{u}_{\mathcal{G}}(\omega,y)>x\}.$
Observe that strict monotonicity and continuity of $\mathbb{u}_{\mathcal{G}}(\omega,\cdot)$ imply $\Phi_{\mathcal{G}}(\omega,\cdot)$ is strictly increasing and continuous over $\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ for all $\omega\in\Omega$. Then, in particular, $\mathbb{u}_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ is a bijection and we can write
$\Phi_{\mathcal{G}}(\omega,x)=\begin{cases}+\infty&\quad\text{for }x\geq\mathbb{u}_{\mathcal{G}}(\omega,+\infty),\\ \mathbb{u}_{\mathcal{G}}^{-1}(\omega,x)&\quad\text{for }x\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}}),\\ -\infty&\quad\text{for }x\leq\mathbb{u}_{\mathcal{G}}(\omega,-\infty).\end{cases}$ (16)
Notice that, fixed $\omega\in\Omega$, whenever $x\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ then $\Phi_{\mathcal{G}}(\omega,x)$ is finite. We claim that $\Phi_{\mathcal{G}}(\cdot,x):\Omega\to\overline{\mathbb{R}}$ is ${\mathcal{G}}$-measurable for all $x\in\mathbb{R}$. To see this, let $x\in\mathbb{R}$ be fixed and consider the set $A=(a,+\infty]\in\mathbb{B}_{\overline{\mathbb{R}}}$ with $a>-\infty$. We show that $\left\{\omega\in\Omega:\Phi_{\mathcal{G}}(\omega,x)\in A\right\}\in{\mathcal{G}}$. We observe that, by strict monotonicity of $\mathbb{u}_{\mathcal{G}}$ the latter coincides with $\left\{\omega\in\Omega:x>\mathbb{u}_{\mathcal{G}}(\omega,a)\right\}$ which is an element of ${\mathcal{G}}$ as $\mathbb{u}_{\mathcal{G}}$ is ${\mathcal{G}}$-measurable.
Addittionally, we wish to emphasize that $\Phi_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\left[-\infty,+\infty\right]$ is continuous for all $\omega\in\Omega$, and therefore $\Phi_{\mathcal{G}}:\Omega\times\mathbb{R}\to\left[-\infty,+\infty\right]$ is a Carathéodory function (see Aliprantis , Definition 4.50). In particular, we observe that $\mathbb{R}$ with the usual euclidean topology is a separable metrizable space and $\overline{\mathbb{R}}=\left[-\infty,+\infty\right]$ with the standard topology on the extended real numbers is metrizable. Then, by virtue of Aliprantis , Lemma 4.51, $\Phi_{\mathcal{G}}$ is ${\mathcal{G}}\otimes\mathbb{B}_{\mathbb{R}}$ measurable.
Remark 5.1
By (16), for any $\omega\in\Omega$ and $x\in\mathrm{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ we have
$\mathbb{u}_{\mathcal{G}}\Big{(}\omega,\Phi_{\mathcal{G}}(\omega,x)\Big{)}=x.$ (17)
Step 4: we conclude proving the desired representation of the conditional Chisini mean. Chosen $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, by strict monotonicity of $\mathbb{u}(\omega,\cdot)$ we have that
$\mathbb{u}(\omega,-\left\|f\right\|_{\infty})\leq\mathbb{u}(\omega,f(\omega))\leq\mathbb{u}(\omega,-\left\|f\right\|_{\infty}),\quad\quad\forall\;\omega\in\Omega.$
Taking a version $g$ of $\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)\middle|{\mathcal{G}}\right]$ we can find $A\in{\mathcal{G}}$ such that $\mathbb{P}(A)=1$ and exploting (15) we may write
$\mathbb{u}_{\mathcal{G}}(\omega,-\left\|f\right\|_{\infty})<g(\omega)<\mathbb{u}_{\mathcal{G}}(\omega,\left\|f\right\|_{\infty})\quad\text{ for all }\omega\in A.$ (18)
Let us define $\tilde{g}=g\mathbf{1}_{A}+0\mathbf{1}_{A^{c}}$ and observe that now $\tilde{g}(\omega)\in\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})$ for all $\omega\in\Omega$. Moreover, $\tilde{g}$ is still a version of the conditional expectation of $f$. Recalling that the mapping $\Phi_{\mathcal{G}}(\omega,\cdot):\mathbb{R}\to\overline{\mathbb{R}}$ is strictly increasing on $\text{Im}_{\omega}(\mathbb{u}_{\mathcal{G}})\subseteq\mathbb{R}$ for all $\omega\in\Omega$, and application to (18) yields
$\Phi_{\mathcal{G}}(\omega,\mathbb{u}_{\mathcal{G}}(\omega,-\left\|f\right\|_{\infty}))<\Phi_{\mathcal{G}}(\omega,\tilde{g}(\omega))<\Phi_{\mathcal{G}}(\omega,\mathbb{u}_{\mathcal{G}}(\omega,\left\|f\right\|_{\infty})).$
Consequently, by (17) $\Phi_{\mathcal{G}}(\omega,\mathbb{u}_{\mathcal{G}}(\omega,x))=x$ for all $\omega\in\Omega$ and all $x\in\mathbb{R}$ and it follows that that $-\left\|f\right\|_{\infty}<\Phi_{\mathcal{G}}(\omega,\tilde{g})<\left\|f\right\|_{\infty}$, for every $\omega\in\Omega$. Hence $\Phi_{\mathcal{G}}\left(\cdot,\tilde{g}\right)\in\mathcal{L}^{\infty}(\Omega,\mathcal{G})$. In particular then we have
$\mathbb{u}_{\mathcal{G}}\Big{(}\omega,\Phi_{\mathcal{G}}\big{(}\omega,\tilde{g}(\omega)\big{)}\Big{)}=\tilde{g}(\omega)\quad\text{ for every }\omega\in\Omega.$ (19)
We prove that $\Phi_{\mathcal{G}}(\omega,\tilde{g}(\omega))$ is the conditional Chisini mean in that it solves the infinite dimensional system of equations given in (5), for the preference representing functional $V_{\Omega}(f)=\int_{\Omega}\mathbb{u}(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ defined in Theorem 3.1. Considering an arbitrary set $A\in{\mathcal{G}}\subseteq{\mathcal{F}}$, we have
$V_{\Omega}\big{(}\Phi_{\mathcal{G}}(\cdot,\tilde{g})\mathbf{1}_{A}\big{)}$ $=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}\big{(}\cdot,\Phi_{\mathcal{G}}\left(\cdot,\tilde{g}\right)\mathbf{1}_{A}\big{)}\right]=\mathbb{E}_{\mathbb{P}}\left[\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}\big{(}\cdot,\Phi_{\mathcal{G}}\left(\cdot,\tilde{g}\right)\big{)}\mathbf{1}_{A}\middle|{\mathcal{G}}\right]\right]$
[…]
=(17)${{=}}\mathbb{E}_{\mathbb{P}}\left[\tilde{g}\mathbf{1}_{A}\right]=\mathbb{E}_{\mathbb{P}}\left[\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)\mathbf{1}_{A}\middle|{\mathcal{G}}\right]\right]=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f)\mathbf{1}_{A}\right]$
$=\mathbb{E}_{\mathbb{P}}\left[\mathbb{u}(\cdot,f\mathbf{1}_{A})\right]=V_{\Omega}(f\mathbf{1}_{A}),$
observing that since $\mathbb{u}(\omega,0)=0$ for all $\omega\in\Omega$ by construction (whence also $\Phi_{\mathcal{G}}(\omega,0)=0$ for all $\omega\in\Omega$) we have
$\begin{cases}\mathbb{u}(\omega,f(\omega))\mathbf{1}_{A}(\omega)=\mathbb{u}\big{(}\omega,f(\omega)\mathbf{1}_{A}(\omega)\big{)}\\ \Phi_{\mathcal{G}}\left(\omega,\tilde{g}(\omega)\right)\mathbf{1}_{A}(\omega)=\Phi_{\mathcal{G}}\big{(}\omega,\tilde{g}(\omega)\mathbf{1}_{A}(\omega)\big{)}\end{cases}\quad\text{ for every }\omega\in\Omega.$
Appendix A Appendix: auxiliary results
Lemma A.1
Let $u:\Omega\times\mathbb{R}\to\mathbb{R}$ and assume that for every $\omega\in\Omega$ the function $x\mapsto u(\omega,x)$ is right continuous with left limit (RCLL), nondecreasing and $u(\omega,0)=0$. Suppose additionally that $u(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$. Then:
  1. (i)
    $u$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}-\mathbb{B}_{\mathbb{R}}$ measurable;
  2. (ii)
    for every $f\in\mathcal{L}^{\infty}(\Omega,\mathcal{F})$, $\omega\mapsto u(\omega,f(\omega))\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$
Proof 5
We prove Item (i) in the spirit of KS91 (Proposition 1.1.13 and Remark 1.1.14) as follows. Let $m>0$. For $n\geq 1$, $k=0,1,...,2\cdot 2^{n}-1$, we define
$u^{(n)}(\omega,s):=u\left(\omega,\frac{\left(k+1\right)m}{2^{n}}-m\right)\;\;\;\;\;\text{ for }\;\;\;k\frac{m}{2^{n}}-m\leq s<\frac{\left(k+1\right)m}{2^{n}}-m,$
and fix
$\begin{cases}u^{(n)}(\omega,s):=u\left(\omega,-m\right),\;\;\;\;\;&\text{ for }s<-m,\\ u^{(n)}(\omega,s):=u\left(\omega,m\right),\;\;\;\;\;&\text{ for }s\geq m.\end{cases}$
[…]
Proposition A.2
Let $u:\Omega\times\mathbb{R}\rightarrow\mathbb{R}$ satisfying
  1. (i)
    for every $\omega\in\Omega$ $u(\omega,\cdot)$ is RCLL, nondecreasing and $u(\omega,0)=0$;
  2. (ii)
    $u(\cdot,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$;
  3. (iii)
    the functional $T_{u}:\mathcal{L}^{\infty}(\Omega,\mathcal{F})\rightarrow\mathbb{R}$ defined by $T_{u}(f):=\int_{\Omega}u(\omega,f(\omega))\mathrm{d}\mathbb{P}(\omega)$ is continuous from below  [^8] For any $(f_{n})_{n}\subset\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})$ such that $f_{n}(\omega)\uparrow_{n}f(\omega)$ for any $\omega\in\Omega$ we have $T_{u}(f_{n})\rightarrow T_{u}(f)$.
Then $u$ is continuous in the following sense:
[…]
Proof 6
In the present proof we shall rely on few concepts related to the theory of stochastic processes and in particular on stopping times. We refer the reader to KS91 , page 6, for an exhaustive treatment of the notions involved. Lemma A.1 guarantees that $T_{u}:\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})\to\mathbb{R}$ is well defined and finite valued. Fix $M>0$ and define $X_{t}(\omega):=u(\omega,t-M)$ for $t\geq 0$, omitting the dependence on $M$ for simplicity. Take also the filtration ${\mathcal{F}}_{t}:={\mathcal{F}},t\geq 0$, which is trivially right continuous by definition: $\bigcap_{s>t}{\mathcal{F}}_{s}={\mathcal{F}}={\mathcal{F}}_{t}$. Then $(X_{t})_{t}$ is an RCLL process, adapted to $({\mathcal{F}}_{t})_{t}$. Set $\Delta X_{0}=0$ and for $t>0$ set $\Delta X_{t}:=X_{t}-\sup_{0\leq s<t}X_{s}=X_{t}-\lim_{s\uparrow t}X_{s}$. Following the idea in So13 , we define $\tau^{\varepsilon}$ as the first jump time for the size $\varepsilon>0$, namely
$\tau^{\varepsilon}:=\inf\{t\geq 0\mid\Delta X_{t}>\varepsilon\}\stackrel{{\scriptstyle(\star)}}{{=}}\inf\{t>0\mid X_{t}-\sup_{0\leq s<t}X_{s}>\varepsilon\},$
where equality ()normal-⋆($\star$)( ⋆ ) follows from posing $\Delta X_{0}=0$. Looking back at $u(\omega,\cdot)$, $\tau^{\varepsilon}(\omega)$ is the first time $u(\omega,\cdot)$ jumps, in the interval $(-M,+\infty)$, with a jump size strictly greater than $\varepsilon$, with the usual convention that such time is set to be +$\infty$+ ∞ if no such jump actually occurs. By So13 , $\tau^{\varepsilon}:\Omega\rightarrow[0,+\infty]$ is a $({\mathcal{F}}_{t})_{t}$-stopping time, hence in particular it is ${\mathcal{F}}$-measurable by definition of such a filtration. Setting
$A_{cont}^{M}:=\{\omega\in\Omega\mid[0,+\infty)\ni t\mapsto X_{t}(\omega)\text{is continuous}\},$
we observe that
$\Omega\setminus A_{cont}^{M}=\bigcup_{n\in\mathbb{N}}\{\tau^{\frac{1}{n}}<+\infty\},$
hence the measurability of $A^{M}_{cont}$ follows. Suppose now that $\mathbb{P}(A^{M}_{cont})<1$, then for some $n$ we would have $\mathbb{P}(\tau^{\frac{1}{n}}<+\infty)>0$ and by $\{\tau^{\frac{1}{n}}<+\infty\}=\bigcup_{N\in\mathbb{N}}\{\tau^{\frac{1}{n}}<N\}$ we could find $N$ big enough such that $\mathbb{P}(\tau^{\frac{1}{n}}\leq N)>0$.
Set, for a fixed pair $n,N$ satisfying $\mathbb{P}(\tau^{\frac{1}{n}}\leq N)>0$, $B:=\{\tau^{\frac{1}{n}}\leq N\}$ and $Z:=\tau^{\frac{1}{n}}\mathbf{1}_{B}\in\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})$. For any fixed $\omega\in B$ we know that $\tau^{\frac{1}{n}}(\omega)=\inf\{t\geq 0\mid\Delta X_{t}(\omega)>1/n\}=\lim_{k\to\infty}t_{k}(\omega)$ for some $\omega$-dependent, non-increasing sequence $t_{k}(\omega)\downarrow_{k}\tau^{\frac{1}{n}}(\omega)$ which satisfies $\Delta X_{t_{k}(\omega)}(\omega)>1/n$ for every $k$. This means that, for any $k$,
$\frac{1}{n}\leq X_{t_{k}(\omega)}(\omega)-\sup_{0\leq s<t_{k}(\omega)}X_{s}(\omega)\leq X_{t_{k}(\omega)}-\sup_{0\leq s<\tau^{\frac{1}{n}}(\omega)}X_{s}(\omega)$
as $\sup_{0\leq s<t_{k}(\omega)}X_{s}(\omega)\geq\sup_{0\leq s<\tau^{\frac{1}{n}}(\omega)}X_{s}(\omega)$. Letting $k\to\infty$ this implies by right continuity of $(X_{t})_{t}$
$\frac{1}{n}\leq X_{\tau^{\frac{1}{n}}(\omega)}(\omega)-\sup_{0\leq s<\tau^{\frac{1}{n}}(\omega)}X_{s}(\omega).$ (20)
Take now $s_{k}:=(Z-\frac{1}{k}\mathbf{1}_{B})^{+}$ for $k\in\mathbb{N}$. $s_{k}$ is ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable and bounded. Observe that since $(X_{t})_{t}$ is RCLL, by KS91 Remark 1.1.14 we have $(t,\omega)\mapsto X_{t}(\omega)$ is $\mathbb{B}_{[0,+\infty)}\otimes{\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable, and $\omega\mapsto(\omega,Y(\omega))$ is ${\mathcal{F}}-{\mathcal{F}}\otimes\mathbb{B}_{[0,+\infty)}$ measurable for every ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable bounded $Y:\Omega\rightarrow[0,+\infty)$. Hence $\omega\mapsto X_{s_{k}(\omega)}(\omega)$ is ${\mathcal{F}}-\mathbb{B}_{[0,+\infty)}$ measurable. By continuity from below, argued at the beginning of this proof,
$T_{u}\big{(}(s_{k}-M)\mathbf{1}_{B}\big{)}\rightarrow_{k}T_{u}\big{(}(Z-M)\mathbf{1}_{B}\big{)}.$ (21)
At the same time, $s_{k}(\omega)=Z(\omega)=0$ for every $\omega\in\Omega\setminus B$ and using (20) $X_{s_{k}(\omega)}(\omega)\leq\sup_{0\leq s<\tau^{\frac{1}{n}}(\omega)}X_{s}(\omega)\leq-\frac{1}{n}+X_{\tau^{\frac{1}{n}}(\omega)}(\omega)$, for any $\omega\in B$. Recalling that $u(\omega,0)=0$ for every $\omega\in\Omega$, we can then write
$T_{u}\big{(}(s_{k}-M)\mathbf{1}_{B}\big{)}$ $=\int_{B}u(\omega,s_{k}(\omega)-M)\mathrm{d}\mathbb{P}(\omega)=\int_{B}X_{s_{k}(\omega)}(\omega)\mathrm{d}\mathbb{P}(\omega)$
(20)${{\leq}}-\frac{1}{n}{\mathbb{P}}(B)+\int_{B}X_{\tau^{\frac{1}{n}}(\omega)}(\omega)\mathrm{d}\mathbb{P}(\omega)$ )
$=-\frac{1}{n}\mathbb{P}(B)+\int_{\Omega}u(\omega,Z(\omega)-M)\mathrm{d}\mathbb{P}(\omega)$
so that $T_{u}((s_{k}-M)\mathbf{1}_{B})+\frac{1}{n}\mathbb{P}(B)\leq T_{u}((Z-M)\mathbf{1}_{B})$ reaching a contradiction with (21), as $\frac{1}{n}\mathbb{P}(B)>0$ does not depend on $k$. Now, we conclude that ${\mathbb{P}}(A^{M}_{cont})=1$ for every arbitrarily fixed $M>0$. Since $A_{cont}=\bigcup_{M\in\mathbb{N}}A^{M}_{cont}$ we conclude that both $A_{cont}\in{\mathcal{F}}$ and ${\mathbb{P}}(A)=1$.
Corollary A.3
Under the same assumptions of Proposition A.2, there exists a function $\widehat{u}$ such that $\widehat{u}(\omega,\cdot)$ is nondecreasing and continuous for every $\omega\in\Omega$, $E=\{\omega\in\Omega\mid u(\omega,x)=\widehat{u}(\omega,x)\,\forall x\in\mathbb{R}\}\in{\mathcal{F}}$ and ${\mathbb{P}}(E)=1$. In particular, $T_{u}(f)=T_{\widehat{u}}(f),\,\forall f\in\mathcal{L}^{\infty}(\Omega,{\mathcal{F}})$. Finally, if $u(\omega,\cdot)$ is strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$, $\widehat{u}(\omega,\cdot)$ can be taken strictly increasing on $\mathbb{R}$ for every $\omega\in\Omega$.
Proof 7
As a consequence of Proposition A.2, $A_{cont}\in{\mathcal{F}}$. Set $A:=A_{cont}$ and define now $\widehat{u}(\omega,x):=u(\omega,x)\mathbf{1}_{A}(\omega)+x\mathbf{1}_{\Omega\setminus A}$. Then clearly $\widehat{u}(\omega,\cdot)$ is continuous and nondecreasing either $\omega\in A$ or $\omega\in\Omega\setminus A$. Further for any fixed $x\in\mathbb{R}$ the map $\omega\mapsto\widehat{u}(\omega,x)$ is ${\mathcal{F}}$ measurable. As $u(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ implies $\widehat{u}(\omega,x)\in{\mathcal{L}}^{1}(\Omega,{\mathcal{F}},\mathbb{P})$ for every $x\in\mathbb{R}$ we can invoke Lemma A.1 to obtain that $\widehat{u}$ is ${\mathcal{F}}\otimes\mathbb{B}_{\mathbb{R}}-\mathbb{B}_{\mathbb{R}}$ measurable, $T_{\widehat{u}}$ is well defined in ${\mathcal{L}}^{\infty}(\Omega,{\mathcal{F}})$ and $T_{u}\equiv T_{\widehat{u}}$.
Ethics declarations
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.